OpenStudy (anonymous):

Help please! thanks! 4x2 + 9y2 = 72 x - y2 = -1 Select all of the following that are solutions to the system shown. so is it (3, 2),(3, -2),(-3, 2),(-3, -2)?

OpenStudy (anonymous):

substitute in each possible solution into both equations and see if they are true still. like this (3, 2) x=3, y=2 4x2 + 9y2 = 72 4(3)^2 + 9(2)^2 = 72 [true] x - y2 = -1 3 - 2^2 = -1 [true] the set (3,2) is a solution repeat this process for the remaining sets